# sample.bib

### v1

@@ Abstract
But with regard to the material world, we can at least go so far as this - we can perceive that events are brought about not by insulated interpositions of Divine power, exerted in each particular case, but by the establishment of general laws.
@@
and Wolfram himself writes [Howwegotherewolfram]:
@@
is rules! And by doing so, it would seem that he completes the removal of anthropocentrism in creation and further divides what Heisenberg calls the Cartesian partition (“God-World-I”) [Heisenberg]. But this assessment is not correct, because the dependence of mathematical truths on the experience of “observers like us”(sic) is a major conclusion of Wolfram’s work [Somehistoricalbackgroundwolfram]:
@@
How far does “our particular perception” allow us to speak on fundamental theories of physics and metaphysics? According to Wolfram, the underlying substrate of reality is the “entangled limit of all possible computations” called “the ruliad”. This conjecture is either as abstract as the unknowable God and unfalsifiable, or so vacuous one might as well say “the universe does things”, or some superposition of the two. Tegmark, on the other hand, asserts reality is all “mathematical structure”—a form of radical platonism that aligns with his acceptance of the Many-Worlds interpretation of quantum mechanics [Tegmark_1998]. Yes, a culmination of phenomena including quantum mechanics, the modern computer, and lack of alternative dogma have enabled what I might describe as gnostic atomism. The following paragraph will explicate a brief history of similar ideas, although a more in-depth history is found in Heisenberg’s book on the philosophy of physics [Heisenberg].
@@
In using the plural “us” and “our” consistently, we can assume Stephen is assuming a common agreement amongst readers, or is employing nosism. Either way, the presumption that many humans have a common sampling of the ruliad is left unexplained, along with why we are and remain localized in space and time. It is unclear how observers are delineated in the ruliad and how this is compatible with the purported independence from the universe—the idea of “observers like us” is simply a brute fact for Wolfram. These proposed restrictions on observers bring to mind the fine-tuning argument for intelligent design: as observers, we happen to be just the right kind to be able to name the underlying structure of reality itself. To quote Wolfram, “There has to be just enough equivalencing” of information by human observers “and not too much.” He admits circularity in his reasoning himself:
@@
So Wolfram’s current universe is dyadic—it consists of the ruliad and observers. If conscious observers were emes, these emes would still follow privileged rules which uphold the aforementioned assumptions regarding the observer. For example, it could be that observers are time-and-space-localized compression emes which sample the ruliad. But Wolfram (and Tegmark) are notably unsatisfied with a privileged rule which runs the universe:
@@
How do we corroborate Stephen’s assertions with Gorard’s definitions and our conscious experience? From the definitions the model of reality is relations between vertices on hypergraphs. A binary relation is defined as a set of ordered pairs of elements from two sets. So from the eternally existing atoms of space spring forth such relations. Already the model seems vacuous because which rules are applied when presumably must evolve according to some hyperrule in the enigmatic and so-called “hyperruliad”—an infinite regress problem arises. Regarding the hyperruliad existing, Wolfram writes the following [Wolfram_ruliad]:
@@
The is reminiscent of the question of the creator of God—“If God exists then who created God?”. Stephen’s rather unsatisfactory answer in the context of the hyperruliad existing is basically “we assert we are not in the hyperruliad because it’s not what we perceive” (presumably speaking on behalf of humankind).
@@
Even granting abstractions without agency like numbers acting as a first cause, I will further show that observer-independent truth is impossible and so claims about external reality are void unless there is an explicit reconciliation. Consider the following objections a proponent of the theory might state:
@@
So what happens if one tries to add a statement that “isn’t true”? The basic answer is that it produces an “explosion” in which the entailment fabric can grow to encompass essentially any statement. From the point of view of underlying rules—or the ruliad—there’s really nothing wrong with this. But the issue is that it’s incompatible with an “observer like us”—or with any realistic idealization of a mathematician.
@@
An appeal to common belief! External intelligence is also proposed in the model, so in the above example Jonathan can be replaced with an alien, or a rock. Is the ruliad no longer an objective concept if an alien mind could not have a conception of this ruliad? To go further, what if our nature as “observers like us” give rise to the abstraction of the ruliad?
@@
But isn’t “formality”, a word also used in mathematics as well as philosophy, laden with an observers presuppositions? Gorard describes Definition 5 as coming from a formalism within mathematical logic, so it appears mathematical logic subsumes graph rewriting systems, and binary relations have an a priori existence undergirding them. Is there a causal chain connecting these concepts and how was this knowledge imprinted onto us particular observers? In Plato’s case, knowledge was recollection from the eternal soul. David Hilbert addressed mathematical theory as relations in 1899 [Blanchette_2007]:
@@
Tegmark’s proof reduces to checking a lookup table of abstract models to physical measurements, which he expresses as a graph isomorphism problem (Fig. 4). This seems rather contrived and not my definition of identically equal: by using the phrase “these two things are identical” one already presupposes distinctness with the word two. Then one has to prove how his or her perception and process of enumeration failed to identify their oneness, rather than claiming they are identically equal because their components can be paired.
@@
I would like for him to think of an example that doesn’t actually have the “substrate out of which the computer is built”, rather than to have to explain this caveat. Then I would be satisfied, knowing that abstractions without agency can act as a first cause.
Wolfram clinically notes that “the only thing special about the universe to us is us ourselves” [Wolfram_toe] and I believe Tegmark, maintaining a great interest in ethics, would agree. Wolfram modifies the Cartesian partition (“God-World-I”) by reintroducing “I” as “observers” and attempts to meld them with “World”. Tegmark removes “I” and “God” and asserts that everything is in some sense “World”. I understand why Heisenberg warned that the Cartesian partition is dangerous. I respect both theories of everything, but find inconsistencies in the reconciliation of “I” the observer and the transcendent truths of mathematics—a cogent narrative has yet to be written which marries potential and instantiated truth. Aristotle thought that potentiality so understood is indefinable, claiming that the general idea can be grasped from a consideration of cases [aristotle]. Like Aristotle, I believe that actuality is to potential as

### v3

@@ Abstract
But with regard to the material world, we can at least go so far as this—we can perceive that events are brought about not by insulated interpositions of Divine power, exerted in each particular case, but by the establishment of general laws.
@@
and Stephen Wolfram himself writes [Howwegotherewolfram]:
@@
is rules! And by doing so, it would seem that he completes the removal of anthropocentrism in creation and further divides what Heisenberg calls the Cartesian partition (“God-World-I”) [Heisenberg]. But this assessment is not correct, because the dependence of mathematical truths on the experience of “observers like us”(sic) is a major conclusion of the theory [Somehistoricalbackgroundwolfram]:
@@
How far does “our particular perception” allow us to speak on fundamental theories of physics and metaphysics? According to Wolfram, the underlying substrate of reality is the “entangled limit of all possible computations” called “the ruliad”. This conjecture is either as abstract as the unknowable God and unfalsifiable, or so vacuous one might as well say “the universe does things”, or some superposition of the two. Max Tegmark, on the other hand, asserts reality is all “mathematical structure”—a form of radical platonism that aligns with his acceptance of the Many-Worlds interpretation of quantum mechanics [Tegmark_1998]. Yes, a culmination of phenomena including quantum mechanics, the modern computer, and lack of alternative dogma have enabled what I might describe as gnostic atomism. The following paragraph will explicate a brief history of similar ideas, although a more in-depth history is found in Heisenberg’s book on the philosophy of physics [Heisenberg].
@@
In using the plural “us” and “our” consistently, we can assume Wolfram is assuming a common agreement amongst readers, or is employing nosism. Either way, the presumption that many humans have a common sampling of the ruliad is left unexplained, along with why we are and remain localized in space and time. It is unclear how observers are delineated in the ruliad and how this is compatible with the purported independence from the universe—the idea of “observers like us” is simply a brute fact for Wolfram. These proposed restrictions on observers bring to mind the fine-tuning argument for intelligent design: as observers, we happen to be just the right kind to be able to name the underlying structure of reality itself. To quote Wolfram, “There has to be just enough equivalencing” of information by human observers “and not too much.” He admits circularity in his reasoning himself:
@@
So Wolfram’s current universe is dyadic—it consists of the ruliad and observers. If conscious observers were emes, these emes would still follow privileged rules which uphold the aforementioned assumptions regarding the observer. For example, it could be that observers are time-and-space-localized compression emes which sample the ruliad. But, in seeming contradiction, Wolfram (and Tegmark) are notably unsatisfied with a privileged rule which runs the universe:
@@
How do we corroborate Wolfram’s assertions with Gorard’s definitions and our conscious experience? From the definitions the model of reality is relations between vertices on hypergraphs. A binary relation is defined as a set of ordered pairs of elements from two sets. So from the eternally existing atoms of space such relations spring forth. Now these relations must be effectively acting at distances like gravity on celestial bodies. Presumably due to their fundamental nature these interactions are inscrutable unless seen by secondary effects. However, once we perform experiments we are thrust into a world of instances, we can weave any narrative we want regarding what we perceive on the detectors, and despite being flexible enough for fitting, the model is vacuous because which rules applied when presumably must evolve according to some hyperrule—the ruliad is a secondary creation. Regarding the so-called “hyperruliad” existing, Wolfram writes the following [Wolfram_ruliad]:
@@
The is reminiscent of the question of the creator of God—“If God exists then who created God?”. Wolfram’s rather unsatisfactory answer in the context of the hyperruliad existing is basically “we assert we are not in the hyperruliad because it’s not what we perceive” (presumably speaking on behalf of humankind).
@@
Even granting abstractions without agency like numbers the ability to create instances, I will further show that observer-independent truth is impossible and so claims about external reality are void unless there is an explicit reconciliation. Consider the following objections a proponent of the theory might state:
@@
So what happens if one tries to add a statement that “isn’t true”? The basic answer is that it produces an “explosion” …From the point of view of underlying rules—or the ruliad—there’s really nothing wrong with this. But the issue is that it’s incompatible with an “observer like us”—or with any realistic idealization of a mathematician.
@@
I don’t believe an appeal to common belief is proof, and this argument relies on the same anthropocentrism that Wolfram and Tegmark desperately try to avoid. External intelligence is also proposed in the model, so in the above example Jonathan can be replaced with an alien, or a rock. Is the ruliad no longer an objective concept if an alien mind could not have a conception of this ruliad? To go further, what if our nature as “observers like us” give rise to the abstraction of the ruliad?
@@
But isn’t “formality”, a word also used in mathematics as well as philosophy, laden with an observers presuppositions? Gorard describes Definition 5 as coming from a formalism within mathematical logic, so it appears mathematical logic subsumes graph rewriting systems, and binary relations have an a priori existence undergirding them. Is there a causal chain connecting these concepts and how was this knowledge imprinted onto us particular observers? In Plato’s case, knowledge of his theory was recollection from the eternal soul. David Hilbert addressed mathematical theory as relations in [Blanchette_2007]:
@@
Tegmark’s proof reduces to checking a lookup table of abstract models to physical measurements, which he expresses as a graph isomorphism problem (Fig. 4). This seems rather contrived and not my definition of identically equal: by using the phrase “these two things are identical” one already presupposes distinctness with the word two. Then one has to prove how his or her perception and process of enumeration failed to identify their oneness, rather than claiming they are identically equal because their respective components can be paired.
@@
I would like for him to think of an example that doesn’t actually have the “substrate out of which the computer is built”, rather than to have to explain this caveat. Then I would be satisfied, knowing that abstractions without agency can act as a first cause and bring about instances.
Wolfram clinically notes that “the only thing special about the universe to us is us ourselves” [Wolfram_toe] and I believe Tegmark, who maintains a great interest in ethics, would agree. Wolfram modifies the Cartesian partition (“God-World-I”) by reintroducing “I” as “observers” and attempts to meld them with “World”. Tegmark removes “I” and “God” and asserts that everything is in some sense “World”. I understand why Heisenberg noted that modern science decided in favor of Plato and warned that the Cartesian partition is dangerous: in both theories of everything I find inconsistencies in the reconciliation of “I” the observer and the transcendent truths of mathematics. A cogent narrative has yet to be written which marries potential and instantiated truth. Aristotle thought that potentiality so understood is indefinable, claiming that the general idea can be grasped from a consideration of cases. Like Aristotle, I believe that actuality is to potential as [aristotle]